# Data Release 1 of the Dark Energy Spectroscopic Instrument

### v1

@@ 4.1 File Access
We anticipate that the majority of users will initiate their data exploration with one of the merged redshift catalogs, spectro/redux/iron/zcatalog/v1/ (containing all objects) or survey/catalogs/dr1/LSS/ (containing objects used in DESI cosmological anlyses), as described in §3.3.2 and §3.4, respectively. Subsequently, users can further refine their data selection by focusing on the objects of particular interest. Once the selection has been defined, users can proceed to download only the files containing spectra relevant to their analysis (see §4.3).
4.2 Database Access
@@
NOIRLab’s Astro Data Lab platform offers anonymous public access343434https://datalab.noirlab.edu/desi to DESI data via a web-query interface and a Table Access Protocol (TAP) handle.353535TAP-aware clients such as TOPCAT (Taylor, 2005) can point to https://datalab.noirlab.edu/tap and select the desi_dr1 database. Additionally, the Astro Data Lab platform offers authenticated access via a JupyterLab363636https://jupyter.org/ server, as well as access to the full-depth DESI spectra through the SPectra Analysis and Retrievable Catalog Lab (SPARCL373737https://astrosparcl.datalab.noirlab.edu), which features a spectral database with a programmatic interface (Juneau et al., 2024b). The subset of spectra available via SPARCL is limited to the 18.7M healpix-coadded spectra (see §3.3.1) that have been combined across cameras. Other types of spectra and files are available at the file-based archive at NERSC previously described in §4.1.
Finally, individuals with NERSC access can query the DESI database using SQL, or they can utilize the pre-installed specprodDB383838https://github.com/desihub/specprod-db Python code, which provides convenient SQLAlchemy (Bayer, 2012) wrapper objects, indexed using q3c (Koposov & Bartunov, 2006, 2019), for rapid, streamlined data-access and manipulation.
4.3 Tutorials
The DESI collaboration creates and maintains a variety of tutorials in the form of Jupyter notebooks (Kluyver et al., 2016; Juneau et al., 2021). These tutorials are intended to facilitate the introduction of various data products and methods for accessing them and are arranged into thematic and topical sub-directories. Due to the significant data volume and complexity, these tutorials may be used as an entry point to get started with DESI data access and analysis. The main GitHub repository for DESI tutorials is located at: https://github.com/desihub/tutorials.
@@ 5 Summary
Finally, we document how the data are organized and can be accessed publicly. The data being released in DR1 include not only individual and coadded spectra and redshift catalogs, but also the LSS catalogs used in all the DESI year-one cosmological analyses and more than 20 VACs spanning a broad range of classes of objects and scientific scope.
All the figures and key statistics in this paper have been produced using DR1 files and Jupyter notebooks and Python code kept in https://github.com/desihub/dr1paper.
@@ Appendix A Guadalupe Spectroscopic Production
DR1 includes Guadalupe as a supplementary spectroscopic production, covering 653 tiles observed in the first two months of the DESI main survey from 2021 May 14 through 2021 July 9. It was run at the same time and using the same software tags as the Fuji production of the SV data which was released in the EDR (see Table 1; DESI Collaboration et al., 2024b). The purpose of Guadalupe was to provide a standardized dataset for early DESI main-survey analyses (see §1.2). We include Guadalupe in DR1 in support of those early publications, but in all other regards it is superseded by the Iron spectroscopic production and is not recommended for any new work.
@@ Appendix B Primary Targets
Main survey bitmasks are recorded in fibermap columns DESI_TARGET, BGS_TARGET, and MWS_TARGET.404040Bitmasks such as SV3_DESI_TARGET (Myers et al., 2023a) are also typically present in the fibermap, but will be zero for Main Survey targets. Similarly, DESI_TARGET will be zero for other survey types. Table 11 lists the DESI_TARGET bits for dark-time targets and calibration targets such as standard stars and sky locations. Table 12 lists the BGS_TARGET bits for the Bright Galaxy Survey. Table 13 lists the MWS_TARGET bits for Milky Way Survey targets, together with some bits that represent calibration targets derived solely from Gaia instead of from Legacy Surveys imaging. These target selection bits are also defined programmatically in the open-source desitarget414141https://github.com/desihub/desitarget software package (Myers et al., 2023b). A YAML-format file describing the bits is in subdirectory py/desitarget/data/targetmask.yaml, with convenience wrapper objects in the Python module desitarget.targetmask. Examples of accessing these bitmasks using this code can be found in §2 of Myers et al. (2023a).
@@
Note. — Bits are stored in the desi_mask and accessed via the DESI_TARGET column (for more details see Myers et al. 2023a).
@@
Note. — Bits are stored in the bgs_mask and accessed via the BGS_TARGET column (see Myers et al. 2023a for more details).
@@
Note. — Bits are stored in the mws_mask and accessed via the MWS_TARGET column (see Myers et al. 2023a for more details).
@@ Appendix C Secondary Targets
In addition to its primary science goals, the DESI survey incorporates a range of “secondary” targets to pursue bespoke research (see §2.2). In this appendix, we describe the secondary target campaigns included in the main survey and outline how the bit-values in their scnd_mask and SCND_TARGET column (see §2.4 of Myers et al., 2023a) can be linked back to the relevant program.424242Bitmasks such as SV3_SCND_TARGET (Myers et al., 2023a) are also typically present in the fibermap, but will be zero for main-survey targets. Similarly, SCND_TARGET will be zero for other survey types.
In Table 14 we list the bit-names and bit-values for secondary targets that were scheduled for observation during the DESI main survey. The vast majority of these secondary targeting programs are documented in Appendix B of DESI Collaboration et al. (2024b). Below, we describe the new programs that were added for the main survey. Further scientific justification for many of these new programs can be found in the MWS overview paper (Cooper et al., 2023). Details about the selection of each type of secondary target are also available at the associated docs link for main survey secondary targets.434343https://data.desi.lbl.gov/public/edr/target/secondary/main/docs
@@
Note. — Bits are stored in the scnd_mask and accessed via the SCND_TARGET column (see Myers et al. 2023a for more details).
@@ C.3 GC_BRIGHT, GC_DARK
Primary target selections typically mask out sources near globular clusters using the Legacy Surveys CLUSTER bit-mask.444444https://www.legacysurvey.org/dr9/bitmasks/ The GC secondary programs are designed to select stars in globular clusters to ameliorate any paucity of DESI targets in these CLUSTER regions. The targets consist of high-probability () cluster members from Vasiliev & Baumgardt (2021). As the GC targets have a very low sky density () they are unlikely to interfere with DESI large scale structure analyses, so they are prioritized just below DESI primary targets and white dwarf targets. The GC_BRIGHT (GC_DARK) sample, intended for observations during bright (dark) time, is limited to (.
@@ C.4 DWF_BRIGHT_HI, DWF_BRIGHT_LO, DWF_DARK_HI, DWF_DARK_LO
A star is considered a “member” of the relevant galaxy if it is within five (three) half-light-radii of the galactic center if the galaxy’s half-light-radius is . The selection is further refined by cross-matching to Gaia EDR3 (Gaia Collaboration et al., 2022) and restricting to likely members based on parallax and proper motion. For stars that are not in Gaia EDR3, imaging from LS/DR9 is used to estimate whether a target is a member of the appropriate galaxy. A star is assigned to be “high” (HI) or “low” (LO) priority based on whether information from Gaia and the Legacy Surveys implies that it is highly likely to be a member of the appropriate dwarf galaxy. We list the resulting numbers of targets and the area covered by each galaxy in Table 15. Finally, targets are assigned to the BRIGHT or DARK class, intended for observations during bright or dark time if they fall in the magnitude range or , respectively.
@@ Appendix D Special Observations
As discussed in §2.3.3, DR1 contains a number of tiles that were dedicated to special observations. These tiles can be identified by having special in the SURVEY column of the DR1 tiles-iron.fits file (see §3.3). The vast majority of science targets for these programs have the SCND_ANY bit set in desi_mask (see Table 11) and the DARK_TOO_LOP or BRIGHT_TOO_LOP bit set in scnd_mask (see Table 14). This is because science targets for rapid-turnaround special tiles are all loosely interpreted to be “Targets of Opportunity.” In this appendix, we briefly describe the DR1 special observations summarized in Table 5.
@@
The special backup tiles were designed to test the backup program target selection algorithms and implementation before enabling routine observations of the backup program in the main survey (Cooper et al. 2023, Dey et al. 2025, in preparation). These tiles may use different versions of targeting than the eventual main survey backup tiles; consequently, we do not recommend combining special and main survey backup tiles.
@@
Tiles 80980 and 80981 were designed to include Targets of Opportunity (see, e.g. §3.2.2 of Myers et al., 2023a) in addition to normal bright targets, in order to test the infrastructure for handling candidate transient sources.
@@
Tile 80977 was a deep tile designed to study the nature of QSO and ELG contaminants in the Sagittarius Stream (e.g. Lynden-Bell & Lynden-Bell, 1995).
@@
The m31 special tiles were designed to study the properties of the halo of M31. These observations are detailed in Dey et al. (2023).
@@
The odin special tile targeted Ly emitters (LAEs) and Lyman-break galaxies (LBGs) selected in the COSMOS field (Scoville et al., 2007). The scientific aim of this program was to investigate the capability of DESI to execute future, higher redshift surveys.
@@
The tertiary1 special tiles obtained spectroscopy for a large fraction of magnitude galaxies in the COSMOS field, in order to provide a very dense and complete set of galaxy redshifts. This program marked the transition from using a variety of names to describe special tiles to incrementing each new special program as a “tertiary” program (see, e.g., Myers et al., 2023a).
@@ Appendix E Known Issues
We did not properly account for a change in the string representation of Boolean values between Gaia DR2 and EDR3 (Lindegren et al., 2018; Gaia Collaboration et al., 2022). Due to this oversight, the Gaia DUPLICATED_SOURCE column is always False for DESI targets that derived their values from EDR3. The implication of this is that no source in DESI targeted using Gaia EDR3 (see §4.1.4 of Myers et al. 2023a) would have been masked on DUPLICATED_SOURCE. Almost 2 million sources in the Iron HEALPix-based catalog were selected using Gaia EDR3. Approximately 2.8% of sources in Gaia EDR3 have the DUPLICATED_SOURCE column set, so roughly 56k of the 2M dark-time sources in DR1 may have been unintentionally targeted.
@@
Redshifts for Ly quasars are systematically underestimated with increasing redshift owing to incorrect modeling of the Lyman series optical depth in Redrock, also present for EDR (Wu & Shen, 2023; Bault et al., 2025). The average bias is modest at , but exceeds 100 km/s at . This will be corrected in future releases, but remains present in Iron. The DR1 ZLyA VAC provides corrected redshifts for Ly quasars.
@@
For most of the tiles in Target Selection Validation (SV1), proper-motion corrections were applied in Fiberassign when the tile was designed.454545The design date can differ from when a tile was observed. A consequence is that the (TARGET_RA, TARGET_DEC, and REF_EPOCH) values are altered to have a REF_EPOCH of the date that the tile was designed, which makes them differ from the input photometric column values. The information is correct and consistent with the photometry, however.
@@ Appendix F Summaries of Value Added Catalogs
This appendix expands upon the set of VACs introduced in §3.5 and summarized in Table 10. Specifically, Appendix F.1 contains general VACs not tied to a specific science case and spanning a variety of scientific applications; Appendix F.2 contains VACs that are primarily concerned with Milky Way or stellar science; Appendix F.3 contains VACs associated with extragalactic science, including voids, galaxy groups, galaxy statistics and strong lenses; Appendix F.4 contains VACs associated with quasar and active galactic nuclei (AGN) science, including summary statistics, updated redshifts, Mg II absorption systems and damped Ly (DLA) systems; and Appendix F.5 contains data products produced by the Ly Forest Working Group which were used in the analysis in DESI Collaboration et al. (2025a).
@@ F.1 General VACs
LS/DR9 Photometry VAC delivers merged targeting catalogs (targetphot) and Tractor464646https://github.com/dstndstn/tractor (Lang et al., 2016) catalog photometry (tractorphot) from the Legacy Surveys DR9 (LS/DR9474747https://www.legacysurvey.org/dr9; Dey et al., 2019) for all observed and potential targets (excluding sky fibers) in DR1.484848https://github.com/moustakas/desi-photometry The observed targets in this VAC correspond to objects with at least one observation in DR1, while the potential targets are the targets that DESI could have observed in a given fiber-assignment configuration (including the objects that were actually observed).

### v2

@@ IV.1 File Access
We anticipate that the majority of users will initiate their data exploration with one of the merged redshift catalogs, spectro/redux/iron/zcatalog/v1/ (containing all objects) or survey/catalogs/dr1/LSS/ (containing objects used in DESI cosmological anlyses), as described in §III.3.2 and §III.4, respectively. Subsequently, users can further refine their data selection by focusing on the objects of particular interest. Once the selection has been defined, users can proceed to download only the files containing spectra relevant to their analysis (see §IV.3).
IV.2 Database Access
@@
NOIRLab’s Astro Data Lab platform offers anonymous public access343434https://datalab.noirlab.edu/desi to DESI data via a web-query interface and a Table Access Protocol (TAP) handle.353535TAP-aware clients such as TOPCAT (taylor05a) can point to https://datalab.noirlab.edu/tap and select the desi_dr1 database. Additionally, the Astro Data Lab platform offers authenticated access via a JupyterLab363636https://jupyter.org/ server, as well as access to the full-depth DESI spectra through the SPectra Analysis and Retrievable Catalog Lab (SPARCL373737https://astrosparcl.datalab.noirlab.edu), which features a spectral database with a programmatic interface (juneau24a). The subset of spectra available via SPARCL is limited to the 18.7M healpix-coadded spectra (see §III.3.1) that have been combined across cameras. Other types of spectra and files are available at the file-based archive at NERSC previously described in §IV.1.
Finally, individuals with NERSC access can query the DESI database using SQL, or they can utilize the pre-installed specprodDB383838https://github.com/desihub/specprod-db Python code, which provides convenient SQLAlchemy (bayer12a) wrapper objects, indexed using q3c (koposov06a; koposov19a), for rapid, streamlined data-access and manipulation.
IV.3 Tutorials
The DESI collaboration creates and maintains a variety of tutorials in the form of Jupyter notebooks (kluyver16a; juneau21a). These tutorials are intended to facilitate the introduction of various data products and methods for accessing them and are arranged into thematic and topical sub-directories. Due to the significant data volume and complexity, these tutorials may be used as an entry point to get started with DESI data access and analysis. The main GitHub repository for DESI tutorials is located at: https://github.com/desihub/tutorials.
@@ V Summary
Finally, we document how the data are organized and can be accessed publicly. The data being released in DR1 include not only individual and coadded spectra and redshift catalogs, but also the LSS catalogs used in all the DESI year-one cosmological analyses and 24 VACs spanning a broad range of astrophysical classes of objects and scientific scope.
All the figures and key statistics in this paper have been produced using DR1 files and Jupyter notebooks and Python code in an open-source repository.404040https://github.com/desihub/dr1paper The Digital Object Identifier (DOI) for DR1 is 10.5281/zenodo.15089588, which includes all the data used to generate the figures and tables in this paper, as well as the figures themselves.414141https://zenodo.org/records/15089588
@@ Appendix A Guadalupe Spectroscopic Production
DR1 includes Guadalupe as a supplementary spectroscopic production, covering 653 tiles observed in the first two months of the DESI main survey from 2021 May 14 through 2021 July 9. It was run at the same time and using the same software tags as the Fuji production of the SV data which was released in the EDR (see Table 1; DESI Collaboration et al., 2024b). The purpose of Guadalupe was to provide a standardized dataset for early DESI main-survey analyses (see §I.2). We include Guadalupe in DR1 in support of those early publications, but in all other regards it is superseded by the Iron spectroscopic production and is not recommended for any new work.
@@ Appendix B Primary Targets
Main survey bitmasks are recorded in fibermap columns DESI_TARGET, BGS_TARGET, and MWS_TARGET.424242Bitmasks such as SV3_DESI_TARGET (Myers et al., 2023) are also typically present in the fibermap, but will be zero for Main Survey targets. Similarly, DESI_TARGET will be zero for other survey types. Table 11 lists the DESI_TARGET bits for dark-time targets and calibration targets such as standard stars and sky locations. Table 12 lists the BGS_TARGET bits for the Bright Galaxy Survey. Table 13 lists the MWS_TARGET bits for Milky Way Survey targets, together with some bits that represent calibration targets derived solely from Gaia instead of from Legacy Surveys imaging. These target selection bits are also defined programmatically in the open-source desitarget434343https://github.com/desihub/desitarget software package (desitarget2023). A YAML-format file describing the bits is in subdirectory py/desitarget/data/targetmask.yaml, with convenience wrapper objects in the Python module desitarget.targetmask. Examples of accessing these bitmasks using this code can be found in §2 of Myers et al. (2023).
@@
Note. — Bits are stored in the desi_mask and accessed via the DESI_TARGET column (for more details see Myers et al. 2023).
@@
Note. — Bits are stored in the bgs_mask and accessed via the BGS_TARGET column (see Myers et al. 2023 for more details).
@@
Note. — Bits are stored in the mws_mask and accessed via the MWS_TARGET column (see Myers et al. 2023 for more details).
@@ Appendix C Secondary Targets
In addition to its primary science goals, the DESI survey incorporates a range of “secondary” targets to pursue bespoke research (see §II.2). In this appendix, we describe the secondary target campaigns included in the main survey and outline how the bit-values in their scnd_mask and SCND_TARGET column (see §2.4 of Myers et al., 2023) can be linked back to the relevant program.444444Bitmasks such as SV3_SCND_TARGET (Myers et al., 2023) are also typically present in the fibermap, but will be zero for main-survey targets. Similarly, SCND_TARGET will be zero for other survey types.
In Table 14 we list the bit-names and bit-values for secondary targets that were scheduled for observation during the DESI main survey. The vast majority of these secondary targeting programs are documented in Appendix B of DESI Collaboration et al. (2024b). Below, we describe the new programs that were added for the main survey. Further scientific justification for many of these new programs can be found in the MWS overview paper (Cooper et al., 2023). Details about the selection of each type of secondary target are also available at the associated docs link for main survey secondary targets.454545https://data.desi.lbl.gov/public/edr/target/secondary/main/docs
@@
Note. — Bits are stored in the scnd_mask and accessed via the SCND_TARGET column (see Myers et al. 2023 for more details).
@@ C.3 GC_BRIGHT, GC_DARK
Primary target selections typically mask out sources near globular clusters using the Legacy Surveys CLUSTER bit-mask.464646https://www.legacysurvey.org/dr9/bitmasks/ The GC secondary programs are designed to select stars in globular clusters to ameliorate any paucity of DESI targets in these CLUSTER regions. The targets consist of high-probability () cluster members from vasiliev21a. As the GC targets have a very low sky density () they are unlikely to interfere with DESI large scale structure analyses, so they are prioritized just below DESI primary targets and white dwarf targets. The GC_BRIGHT (GC_DARK) sample, intended for observations during bright (dark) time, is limited to (.
@@ C.4 DWF_BRIGHT_HI, DWF_BRIGHT_LO, DWF_DARK_HI, DWF_DARK_LO
A star is considered a “member” of the relevant galaxy if it is within five (three) half-light-radii of the galactic center if the galaxy’s half-light-radius is . The selection is further refined by cross-matching to Gaia EDR3 (gaia-collaboration22a) and restricting to likely members based on parallax and proper motion. For stars that are not in Gaia EDR3, imaging from LS/DR9 is used to estimate whether a target is a member of the appropriate galaxy. A star is assigned to be “high” (HI) or “low” (LO) priority based on whether information from Gaia and the Legacy Surveys implies that it is highly likely to be a member of the appropriate dwarf galaxy. We list the resulting numbers of targets and the area covered by each galaxy in Table 15. Finally, targets are assigned to the BRIGHT or DARK class, intended for observations during bright or dark time if they fall in the magnitude range or , respectively.
@@ Appendix D Special Observations
As discussed in §II.3.3, DR1 contains a number of tiles that were dedicated to special observations. These tiles can be identified by having special in the SURVEY column of the DR1 tiles-iron.fits file (see §III.3). The vast majority of science targets for these programs have the SCND_ANY bit set in desi_mask (see Table 11) and the DARK_TOO_LOP or BRIGHT_TOO_LOP bit set in scnd_mask (see Table 14). This is because science targets for rapid-turnaround special tiles are all loosely interpreted to be “Targets of Opportunity.” In this appendix, we briefly describe the DR1 special observations summarized in Table 5.
@@
The special backup tiles were designed to test the backup program target selection algorithms and implementation before enabling routine observations of the backup program in the main survey (Cooper et al., 2023; dey25a)). These tiles may use different versions of targeting than the eventual main survey backup tiles; consequently, we do not recommend combining special and main survey backup tiles.
@@
Tiles 80980 and 80981 were designed to include Targets of Opportunity (see, e.g. §3.2.2 of Myers et al., 2023) in addition to normal bright targets, in order to test the infrastructure for handling candidate transient sources.
@@
Tile 80977 was a deep tile designed to study the nature of QSO and ELG contaminants in the Sagittarius Stream (e.g. lynden-bell95a).
@@
The m31 special tiles were designed to study the properties of the halo of M31. These observations are detailed in dey23a.
@@
The odin special tile targeted Ly emitters (LAEs) and Lyman-break galaxies (LBGs) selected in the COSMOS field (scoville07a). The scientific aim of this program was to investigate the capability of DESI to execute future, higher redshift surveys.
@@
The tertiary1 special tiles obtained spectroscopy for a large fraction of magnitude galaxies in the COSMOS field, in order to provide a very dense and complete set of galaxy redshifts. This program marked the transition from using a variety of names to describe special tiles to incrementing each new special program as a “tertiary” program (see, e.g., Myers et al., 2023).
@@ Appendix E Known Issues
We did not properly account for a change in the string representation of Boolean values between Gaia DR2 and EDR3 (lindegren18a; gaia-collaboration22a). Due to this oversight, the Gaia DUPLICATED_SOURCE column is always False for DESI targets that derived their values from EDR3. The implication of this is that no source in DESI targeted using Gaia EDR3 (see §4.1.4 of Myers et al. 2023) would have been masked on DUPLICATED_SOURCE. Almost 2 million sources in the Iron HEALPix-based catalog were selected using Gaia EDR3. Approximately 2.8% of sources in Gaia EDR3 have the DUPLICATED_SOURCE column set, so roughly 56k of the 2M dark-time sources in DR1 may have been unintentionally targeted.
@@
Redshifts for Ly quasars are systematically underestimated with increasing redshift owing to incorrect modeling of the Lyman series optical depth in Redrock, also present for EDR (wu23a; Bault et al., 2025). The average bias is modest at , but exceeds 100 km/s at . This will be corrected in future releases, but remains present in Iron. The DR1 ZLyA VAC provides corrected redshifts for Ly quasars.
@@
For most of the tiles in Target Selection Validation (SV1), proper-motion corrections were applied in Fiberassign when the tile was designed.474747The design date can differ from when a tile was observed. A consequence is that the (TARGET_RA, TARGET_DEC, and REF_EPOCH) values are altered to have a REF_EPOCH of the date that the tile was designed, which makes them differ from the input photometric column values. The information is correct and consistent with the photometry, however.
@@ Appendix F Summaries of Value Added Catalogs
This appendix expands upon the set of VACs introduced in §III.5 and summarized in Table 10. Specifically, Appendix F.1 contains general VACs not tied to a specific science case and spanning a variety of scientific applications; Appendix F.2 contains VACs that are primarily concerned with Milky Way or stellar science; Appendix F.3 contains VACs associated with extragalactic science, including voids, galaxy groups, galaxy statistics and strong lenses; Appendix F.4 contains VACs associated with quasar and active galactic nuclei (AGN) science, including summary statistics, updated redshifts, Mg II absorption systems and damped Ly (DLA) systems; and Appendix F.5 contains data products produced by the Ly Forest Working Group which were used in the analysis in DESI Collaboration et al. (2025b).
Some of the VACs described below are finalized and will not change (e.g., those with published companion papers), while others associated with papers currently in preparation may be updated after the public data release. The most current version of each VAC and its associated documentation can always be found at the DESI VAC portal.484848https://data.desi.lbl.gov/doc/vac
@@ F.1 General VACs
LS/DR9 Photometry VAC delivers merged targeting catalogs (targetphot) and Tractor494949https://github.com/dstndstn/tractor (lang16a) catalog photometry (tractorphot) from the Legacy Surveys DR9 (LS/DR9505050https://www.legacysurvey.org/dr9; Dey et al., 2019) for all observed and potential targets (excluding sky fibers) in DR1.515151https://github.com/moustakas/desi-photometry The observed targets in this VAC correspond to objects with at least one observation in DR1, while the potential targets are the targets that DESI could have observed in a given fiber-assignment configuration (including the objects that were actually observed).
@@
BAO Cosmology Results VAC provides the cosmology chains and posterior maximization results derived from the BAO analysis in DESI Collaboration et al. (2025a).
Full Shape Cosmology Results VAC provides the cosmology chains and posterior maximization results used full shape analysis in DESI Collaboration et al. (2024c).